# Can Large Language Models Capture Video Game Engagement?

### v1

@@ IV-A2 GPT-4o
GPT4 is, at the time of writing, the most recent of a series of Generative Pre-trained Transformer (GPT) models developed by OpenAI. GPT4 is a closed source model. While a technical report about GPT4 has been published [20], the exact architecture and training data is unknown. What is known is that GPT4 uses a transformer architecture for both vision and language tasks, relies on reinforcement learning from human feedback and makes use of rule-based reward models based on hidden policy models and human-written rubrics to steer the algorithm in a direction that is considered “safe“ by OpenAI. In this paper we use the GPT-4o (Omni) 2024-08-06 model variant. At the time of writing this is considered the flagship model of OpenAI. Unlike previous iterations, GPT-4o is trained end-to-end to incorporate text, audio, image, and video in both its input and output space [60]. We have selected this model because it is one of the most popular [61], state-of-art, closed-source LLMs as an alternative to the open-source LLaVA. We leverage the Open AI API [^6] https://platform.openai.com/ for all reported experiments with GPT-4o.
IV-B Multimodal Input
In our experiments with Multimodal Input, we feed the models with both visual input and a corresponding text prompt. To provide the visual input we first extract single frames from GameVibe videos at a given interval. Then each frame is cropped to a square and downscaled to a fixed size. Particularly, in our experiments using one image we downscale our images to $336\times 336$ pixels to be able to achieve the highest resolution input possible when combining two images in LLaVA models. [^7] LLaVA models support $672\times 672$, $336\times 1344$, $1344\times 336$ resolutions.
More information: https://ollama.com/library/llava
In our early experiments with Multimodal Input, we use a single image as the model’s input due to a limitation of the LLaVA models, which can only consider one image at a time. To circumvent this limitation we stitch the two video frames together vertically (i.e. a top and a bottom image), leaving a white band of $50$ pixels between them. We call this experimental setting Multimodal Input - 1 Image (Stitched). This type of image stitching performs well on LLaVA models compared to other approaches—such as concatenating the visual tokens [62]. For consistency we follow the same processing method with our GPT-4o models when it comes to experiments using a single image. We show an example of this prompting strategy and the output it produces in the Appendix (see Fig. 10).
In experiments involving few-shot prompting, we use two separate images per prompt. This experimental setting, named Multimodal Input - 2 Images, is only applicable to GPT-4o. This choice is partly informed by the aforementioned technical limitation of LLaVA since the few-shot experiments require multiple prompts with multiple images to be chained together. Since these experiments run exclusively on GPT-4o models we downscale images to $512\times 512$ pixels in an effort to exploit the larger input space of GPT vision models [^8] GPT vision models process images in $512\times 512$ pixel tiles with a maximum image size of $2048\times 768$.
More information: https://platform.openai.com/docs/guides/vision
IV-C Text Input
@@
In our experiments using Text Input, we feed the models with text descriptions of two video frames as part of the prompt. We obtain these descriptions using the same LLM we use to generate the engagement evaluation. Similarly to the Multimodal Input - 1 Image setup, we downsample the obtained video frames to $336\times 336$ pixels. Contrary to the previous setup, here we use these images one-by-one and generate descriptions in two different ways. We call these Basic and Advanced Descriptions based on the amount of context given to the model. For the former, we instruct the model to give a brief description, capturing only essential details without subjective commentary based on the setting and layout, enemies, and player action. For obtaining Advanced Descriptions, we instruct the model to also take player engagement into account and generate a description that captures how it might engage the player or viewer. We illustrate this process in the Appendix; see Figs. 11 and 12 respectively. For the engagement prediction task, we feed these descriptions to the models in pairs as part of their text prompt. We show an example of this prompting strategy and the output it produces in the Appendix (see Fig. 13).
IV-D Prompting Methods
@@
All prompting strategies we use for the engagement evaluation task follow a Chain-of-Thought (CoT) paradigm [63, 64]. We ask the models to provide a comparison between the given input frames, reasoning its analysis of engagement, and finally offering a one-word decision (i.e., engagement increase or decrease). Additionally, for the Multimodal Input experiments we also generate a description of the visual input before the comparison. In the Multimodal Input - 1 Image and Text Input experiments the decision is to pick the most engaging frame (see Fig. 10 in the Appendix). In the Multimodal Input - 2 Images experiments, instead, we refine the prompt and ask the model to explicitly output increasing and decreasing labels. We instruct the model to output its answers in a JSON format, which we parse and extract the final decision from; see also Fig. 14 in the Appendix.
@@
For our few-shot experiments in the Multimodal Input - 2 Images setup, we generate artificial reasoning samples for a positive and negative example for each task. We use the same CoT prompt for this process as for the one-shot Multimodal Input - 2 Images experiments. We will call this prompt “CoT prompt” in the remainder of this section. To generate these samples we take the following steps (see also Multimodal Input - 2 Images, Few-Shot in the middle of Fig 2):
@@
    We take a random example from the same game as presented in the task from an unseen session.
@@
By removing incorrect options but using the same CoT prompt when generating positive and negative examples, we ensure that the algorithm’s output is formatted the same way as for the downstream task, including the description, comparison, reasoning, and decision. We use these outputs to construct an artificial history of positive and negative examples, which are added to the final prompt for the engagement evaluation task. For this final step we provide the CoT prompt with the example images as a question, and the example output as an answer; then finally we provide a set of unseen images with the CoT prompt and instruct the LLM to evaluate engagement the same way it would for a one-shot experiment. Figures 15 and 16 in the Appendix detail the process starting from example generation all the way to engagement prediction.
@@ V Results
Figure 4: Application of the temporal shift ($\Delta t$) hyperparameter to the ground truth. The top red bar (Vision Input) shows an example of individual frames extracted from the gameplay video at a 3-second interval. The bottom green bar (Ground Truth) shows a $\Delta t$ of $-2$ seconds, which means that each window aggregates information $2$ seconds before and $1$ second after the corresponding video frame.
Figure 5: Sensitivity analysis across hyperparameters $\Delta t$ and $\theta$. The table presents $\Delta A$ values (relative gain in accuracy). $\Delta t$ is the relative shift of the time window to the frame, and $\theta$ is the binary threshold for the split criterion (i.e., increasing or decreasing engagement). The last column shows average $\Delta A$ across all games.
@@ V-A Experimental Setup
    A preference threshold ($\theta$), taking values between 00 and $1$, that determines whether a difference between the ground truth value of two consecutive time windows is considered a change (increase or decrease) in engagement; e.g. $\theta=0.05$ considers windows which have a difference of more than $5\%$ when evaluating engagement change.
@@
We formulate the downstream task of LLMs as binary classification, and ask our models to predict the increase or decrease of perceived engagement between two frames of consecutive time-windows. We discard predictions which could not be interpreted when either for the following occurs: a) the algorithm predicts no change in engagement, b) the LLM generates outputs we could not parse, or c) the model is not able to provide an output. [^9] Good examples of these cases were LLaVA models providing verbose answers instead of picking one of the provided options for their final answer, e.g.: “It depends on personal preference. If one prefers an immersive experience similar to the player’s perspective, the upper picture might be considered more engaging. On the other hand, if one values breadth and variety in game views, the lower image could be seen as a more engaging alternative.” instead of simply “top” or “bottom” when picking which image is more engaging; and GPT-4o refusing to provide analysis, e.g.: “I’m unable to analyse the content of these images. If you can describe the frames, I can help evaluate the change in engagement”.
We define our baseline performance as equivalent to always predicting the majority class of a given game session. It is worth noting that the baseline differs widely across games and varies to a lesser degree based on the hyperparameter values selected; the lowest baseline across game sessions, on average, is $52\%$ in Far Cry (2004), Blitz Brigade (2013) and Fortnite (2018) whereas the highest is $78\%$ in Counter Strike 1.6 (2003), followed by Heretic (1994) with $73\%$, and Corridor 7 (1994) with $70\%$. Given this level of discrepancy among baselines, and to be able to meaningfully compare performance different instructions [65, 6], we report accuracy gain over the corresponding baseline—instead of the accuracy values per se—as follows:
$\Delta A=\frac{A_{LLM}-A_{b}}{A_{b}}$ (1)
@@
where $\Delta A$ is the relative gain in accuracy; $A_{LLM}$ is the accuracy of the given LLM; and $A_{b}$ is the baseline accuracy given by the majority class. We use the $\Delta A$ measure of performance in all reported experiments in this paper.
V-B Sensitivity Analysis
@@
We experiment with the temporal shift $\Delta t\in\{0,-0.5,-1,-1.5,-2,-2.5,-3\}$ and preference threshold $\theta\in\{0,0.01,0.05,0.1\}$ parameters—introduced in the previous section—using the $7$, $13$, and $34$ billion parameter version of LLaVA, and GPT-4o. The combinations of these parameters, however, result in $112$ experimental setups for each game. Due to space considerations we only present the best performing subset of these hyperparameters ($\Delta t\in\{-1,-2\}$ and $\theta\in\{0.01,0.05\}$). We run these experiments with the Multimodal Input - 1 Image strategy as described in Section IV-B. We chose this setup for the initial parameter tuning because this is the most straightforward setup involving only one image and one text prompt.
Figure 5 presents the $\Delta A$ performance across two $\Delta t$ and $\theta$ values. We can observe that larger $\Delta t$ and $\theta$ values tend to yield higher performance; it also appears that the model size and architecture have a higher impact on $\Delta A$. While LLaVA-7b and LLaVA-34b consistently perform significantly worse than the baseline—measured with Student’s t-Test at significance level $\alpha<0.05$ corrected with the Bonferroni method, accounting for repeated measurements—GPT-4o shows performance comparable to the baseline. Interestingly, LLaVA-13b outperforms the larger LLaVA model and is not significantly worse than the baseline performance.
The best performing hyperparameter set is $\Delta t=-2$ and $\theta=0.05$ both in terms of average and single-game performance. The best performances are as follows: LLaVa-34b improves the baseline by $37\%$ on Blitz Brigade; and GPT-4o by $33\%$, $29\%$, and $26\%$ on Doom (1993), Wolfram (2012), and Blitz Brigade, respectively. Interestingly, we can see comparable performances with other models and configurations on single games. The most indicative of these is the LLaVA-7b model reaching $37\%$ higher performance than the baseline on Wolfram with $\Delta t=-1$ and $\theta=0.01$. The average performance of the aforementioned setup, however, is lower that the performance of models tuned to $\Delta t=-2$ and $\theta=0.05$. This indicates that the models are sensitive to the games themselves and can’t perform uniformly well across the whole dataset. Two striking examples are LLaVA-13b, consistently outperforming every other model on Void Bastards (2019) and LLaVA-7b, consistently underperforming on CS:GO - Dust2 (2012).
Some games are easier to predict than others, regardless of experimental setup. For example, Wolfram, Blitz Brigade, and PUBG are constantly listed within the top performing games in terms of $\Delta A$, whereas Heretic, Counter Strike 1.6, Overwatch 2 (2022), and HROT (2023) yield among the lowest $\Delta A$. It is important to note that games where engagement changes are predicted well by LLMs tend to have lower baselines (i.e. Wolfram: $57\%$; Blitz Brigade $52\%$; PUBG: $60\%$) whereas games where engagement is not predicted as well tend to have high baselines (i.e. Heretic: $73\%$; Counter Strike 1.6: $78\%$; Overwatch 2: $69\%$; and HROT: $61\%$). This indicates that engagement prediction is easier in game videos that feature more dynamic gameplay footage and a more uniform distribution of increasing vs. decreasing engagement labels.
@@
Considering the overall performance of LLM engagement prediction across games, we fix our parameters for processing the ground truth at $\Delta t=-2s$ and $\theta=0.05$ for the remaining experiments presented in this paper. As we observed high levels of performance across different LLM models, we continue our investigations experimenting with both LLaVA and GPT-4o models.
V-C Text-based Engagement Prediction
Figure 6: $\Delta A$ (relative gain in accuracy) values across games of models fed with Text Input compared to the best LLMs using Multimodal Input - 1 Image presented in Section V-B. Experiments using Text Input are based on text-descriptions of frames only. Experiments using Multimodal Input use both images and text prompts. The rightmost column shows $\Delta A$ values averaged across all games.
@@ V-C Text-based Engagement Prediction
In this section we examine the impact of text-based vs. multimodal prompting strategies on LLM performance. While in the former case we provide solely a text prompt to the model, in the latter case we feed both a text prompt and a corresponding image. Because the performance of LLMs can be affected even by small prompt variations [10], we experiment with both Basic and Advanced prompts. The prompting procedure for the text-based experiments are detailed in Section IV-C. Figure 6 shows the $\Delta A$ performance of Text Input experiments compared to the best Multimodal Input - 1 Image models discussed in the previous section.
@@
In this section our analysis focuses on the Text Input compared to the Multimodal - 1 Image (Stitched) results presented in the previous section. This focus on text allows us to compare the Text Input method to a simple multimodal approach across different models. Our hypothesis is that the strategy of generating text-descriptions of frames first and then using these descriptions as part of the Text Input will improve model performance, because it essentially encodes the images in terms of action and player involvement. We thus assume that using this type of Text Input will present a better representation by discarding surface-level differences between frames and emphasising the structural differences.
Overall, we can note that LLaVA-34b models perform significantly worse than the baseline across all modalities (Multimodal and Text) except when the text-only input is combined with Advanced Descriptions, but the performance still remains on the lower end of the spectrum. LLaVA-13b models yield performance values that are significantly below the baseline interdependently of the description setup. Finally, LLaVA-7b underperforms significantly on the multimodal task. It is somewhat surprising that while the larger LLaVA models generally perform better on multimodal tasks, the smallest model (7b) marginally outperforms the other two larger models of the LLaVA family when fed with text-only input. We hypothesise that this is due to the larger models’ stronger tendency to fall into what Chochlakis et al. [16] call “gravity wells of knowledge priors”. This hypothesis is reinforced when we look at the best performing LLaVA models of Fig. 6. The better performing LLMs are usually fed with Basic instead of Advanced Descriptions. The added context seems to confuse the LLM or fails to orient the models to make accurate predictions. The same issue doesn’t seem to affect the GPT-4o model which performs consistently close to the baseline and better than the LLaVA family overall. While the GPT-4o model performs marginally better on the text-input task using the Advanced Descriptions, the biggest improvement can be observed with Basic Descriptions on Doom with a $39\%$ relative gain in accuracy.
[…]

### v2

@@ IV-B6 GPT-4o
GPT4 is, at the time of writing, the most recent of a series of Generative Pre-trained Transformer (GPT) models developed by OpenAI. GPT4 is a closed source model. While a technical report about GPT4 has been published [51], the exact architecture and training data is unknown. What is known is that GPT4 uses a transformer architecture for both vision and language tasks, relies on reinforcement learning from human feedback and makes use of rule-based reward models based on hidden policy models and human-written rubrics to steer the algorithm in a direction that is considered “safe” by OpenAI. In this paper we use the GPT-4o (Omni) 2024-08-06 model variant. At the time of writing this is considered the flagship model of OpenAI. Unlike previous iterations, GPT-4o is trained end-to-end to incorporate text, audio, image, and video in both its input and output space [52]. We have selected this model because it is one of the most popular [65], state-of-art, closed-source LLMs as an alternative to the open-source LLaVA. We leverage the Open AI API [^11] https://platform.openai.com/ for all reported experiments with GPT-4o.
IV-C Multimodal Input
In our Multimodal experiments we feed the models with both visual input and a text prompt. We extract frames from the GameVibe dataset at a given interval. Despite our selected models capable of handling varying resolutions, for consistency, each frame is cropped to a square and downsized to $512\times 512$. We chose this size because GPT-4o processes images in $512\times 512$ pixel tiles with a maximum image size of $2048\times 768$ [^12] See more information in the OpenAI cookbook:
https://platform.openai.com/docs/guides/vision
. In our Multimodal One-Shot experiments we provide two separate images (consecutive frames) to the model. In out Multimodal Few-Shot experiments, we generate the positive and negative examples in a similar way, using two images and a text prompt as input. For the downstream task of engagement evaluation we provide $6$ images in total for every query (this includes both examples and the final query).
IV-D Text Input
@@
In our experiments using Text Descriptions we feed the models with text descriptions of two video frames as part of the prompt. We obtain these descriptions using the same LLM we use to generate the engagement evaluation. The image input of this task is formatted in the same way as for the Multimodal One-Shot experiments. Contrary to these experiments, however, here we use these images one-by-one and generate descriptions in two different ways. We call these Basic and Advanced Descriptions based on the amount of context given to the model. For the former, we instruct the model to give a brief description, capturing only essential details without subjective commentary based on the setting and layout, enemies, and player action. For obtaining Advanced Descriptions, we instruct the model to also take player engagement into account and generate a description that captures how it might engage the player or viewer. An illustration this process is given in the Appendix; see Figs. LABEL:annex:text-basic and LABEL:annex:text-advanced respectively. For the engagement prediction task, we feed these descriptions to the models in pairs as part of their text prompt. An example of this prompting strategy and the output it produces is given in the Appendix (see Fig. LABEL:annex:text-prediction).
IV-E Prompting Methods
@@
All prompting strategies we use for the engagement evaluation task follow a Chain-of-Thought (CoT) paradigm [78, 90]. We ask the models to provide a comparison between the given input frames, reasoning its analysis of engagement, and finally offering a one-word decision (i.e., engagement increase or decrease). Additionally, for the Multimodal experiments we also generate a description of the visual input before the comparison. In the Multimodal experiments the task of the model is to explicitly output increasing and decreasing labels. We instruct the model to output its answers in a JSON format, which we parse and extract the final decision from; see also Fig. LABEL:annex:oneshot given in the Appendix. In the experiments with Text Descriptions the decision is to pick the most engaging frame (see Fig. LABEL:annex:text-prediction given in the Appendix).
@@
For our Multimodal Few-Shot experiments, we generate artificial reasoning samples for a positive and negative example for each task. We use the same CoT prompt for this process as for the One-Shot experiments. We will call this prompt “CoT prompt” in the remainder of this section. To generate these samples we take the following steps (see also Multimodal Few-Shot in the middle of Fig 2):
@@
    We take a random example from the same game as presented in the task from an unseen play session.
@@
By removing incorrect options but using the same CoT prompt when generating positive and negative examples, we ensure that the algorithm’s output is formatted the same way as for the downstream task, including the description, comparison, reasoning, and decision. We use these outputs to construct an artificial history of positive and negative examples, which are added to the final prompt for the engagement evaluation task. For this final step we provide the CoT prompt with the example images as a question, and the example output as an answer; then finally we provide a set of unseen images with the CoT prompt and instruct the LLM to evaluate engagement the same way it would for a one-shot experiment. Figures LABEL:annex:fewshot-examples and LABEL:annex:fewshot-prediction given in the Appendix detail the process starting from example generation all the way to engagement prediction.
@@ V Results
Figure 4: Application of the temporal shift ($\Delta t$) hyperparameter to the ground truth. The top red bar (Vision Input) shows an example of individual frames extracted from the gameplay video at a 3-second interval. The bottom green bar (Ground Truth) shows a $\Delta t$ of 2-2 seconds, which means that each window aggregates information 22 seconds before and 11 second after the corresponding video frame.
Figure 5: Sensitivity analysis across hyperparameters $\Delta t$ and $\theta$. The table presents $\Delta A$ values (relative gain in accuracy). $\Delta t$ is the relative shift of the time window to the frame, and $\theta$ is the binary threshold for the split criterion (i.e., increasing or decreasing engagement). The last column shows average $\Delta A$ across all games. The last two rows in each experimental setup ($\Delta t$ and $\theta$ combination) show the MLP and CNN baselines.
@@ V-A Experimental Setup
    A preference threshold ($\theta$), taking values between 0 and $1$, that determines whether a difference between the ground truth value of two consecutive time windows is considered a change (increase or decrease) in engagement; e.g. $\theta=0.05$ considers windows which have a difference of more than $5\%$ when evaluating engagement change. We discard labels that vary less than $\theta$, under the assumption that the change—and therefore the decision—is negligible.
@@
While the evaluation follows an ordinal paradigm, we formulate the downstream task of LLMs as binary classification, and ask our models to predict the increase or decrease of perceived engagement between two frames of consecutive time-windows. We discard predictions which could not be interpreted when either for the following occurs: a) the algorithm predicts no change in engagement, b) the LLM generates outputs we could not parse, or c) the model is not able to provide an output. [^13] Good examples of these cases are models not following formatting instructions; providing other answers than “increase” and “decrease” (e.g. “same”); outputting repeating tokens; and—mostly in case of GPT-4o—refusing to provide analysis for unknown reasons, e.g.: “I’m unable to analyse the content of these images. If you can describe the frames, I can help evaluate the change in engagement”.
Based on how the intensity of the action changes in the footage, the class distribution of increasing and decreasing labels can vary between games. At the lower end of the class imbalance a naive Zero Rule Classifier (ZeroR), always predicting the majority class, would have $52\%$ accuracy (e.g. in Far Cry (2004), Blitz Brigade (2013) and Fortnite (2018)) and at the higher end it would have between $70\%$ to $78\%$ (e.g $78\%$ in Counter Strike 1.6 (2003), $73\%$ in Heretic (1994), and $70\%$ in Corridor 7 (1994). To be able to meaningfully compare performance across different games and instructions, we report accuracy gain over a ZeroR baseline [2, 80]—instead of the accuracy values per se—as follows:
$\Delta A=\frac{A_{m}-A_{0}}{A_{0}}$ (1)
@@
where $\Delta A$ is the relative gain in accuracy; $A_{m}$ is the accuracy of the given model; and $A_{0}$ is the accuracy of a ZeroR classifier for a given game. We use the $\Delta A$ measure of performance in all reported experiments in this paper.
V-B Multimodal One-Shot Experiments and Sensitivity Analysis
@@
We experiment with the temporal shift $\Delta t\in\{0,-0.5,-1,-1.5,-2,-2.5,-3\}$ and preference threshold $\theta\in\{0,0.01,0.05,0.1\}$ parameters—introduced in the previous section. Because we experiment on $7$ different algorithms including the MLP and CNN Baseline the combinations of these parameters result in $196$ experimental setups for each game. Due to space considerations we only present the best performing subset of these hyperparameters ($\Delta t\in\{-1,-2\}$ and $\theta\in\{0.01,0.05\}$). We run these experiments with the Multimodal One-Shot strategy as described in Section IV-C. We chose this setup for the initial parameter tuning because this is the most straightforward setup.
Figure 6 presents the $\Delta A$ performance across two $\Delta t$ and $\theta$ values. We can observe that larger $\Delta t$ and $\theta$ values tend to yield higher performance; and it also appears that model size has a high impact on $\Delta A$. Across all experimental setups, we see Phi 3.5 with 4.15b parameters performing the worst of the open source models, while the best models have 8b-8.3b parameters (LLaVA-OV and Qwen2.5 VL, respectively). To measure the significance of the difference between the models and experimental setups we employ repeated paired Student’s t-Tests. We apply the Benjamini-Yekutieli Method [10] to correct our p-values for multiple observations and consider a significance level of $\alpha<0.05$. We make three types of comparisons: 1) we compare models within each experimental configuration ($\Delta t$ and $\theta$ combinations); 2) we compare different experimental configurations within each model; and 3) we compare performance on individual games across all LLMs (but notbaseline models) within each experimental configuration. In the first two cases, our degree of freedom is $19$, while in the last case it is $5$. We make $28$, 66, and $190$ comparisons per test respectively.
Despite GPT-4o outperforming all other algorithms at every turn, after the correction we find no significant difference between models or different experimental setups. Nevertheless the best performing hyperparameter set is $\Delta t=-2$ and $\theta=0.05$ in terms of average performance. Focusing on this setup, we observe that the best overall performance can be attributed to GPT-4o with $1\%$ $\Delta A$; while the worst LLM is Phi 3.5 with an overall $13\%$ $\Delta A$. When it comes to the comparison against the baselines, all LLMs outperform the MLP Baseline across all experimental setups; and generally provide comparable performance to the CNN Baseline.
Interestingly, while GPT-4o provides the best performance, we can see comparable results with other models and configurations on single games. Most notably Doom (1993) is remarkably easy to predict for all models, including the MLP Baseline, while Counter Strike 1.6 (2003) is the hardest. The difference in performance is significant, $t(5)=27.2$, $p=0.001$ (corrected). In fact, Doom is significantly easier to predict than Corridor 7 (1994), $t(5)=9.6$, $p=0.023$; HROT (2023), $t(5)=8.7$, $p=0.026$; and Team Fortress 2 (2007), $t(5)=6.9$, $p=0.048$ —and it is not the only one. Wolfram (2012) yields significantly better results than Corridor 7, $t(5)=13.2$, $p=0.012$; Medal of Honor (2010), $t(5)=13.8$, $p=0.012$; Counter Strike 1.6, $t(5)=10.1$, $p=0.023$; PUBG (2018), $t(5)=9.8$, $p=0.023$; and HROT, $t(5)=9.3$, $p=0.023$; and Blitz Brigade (2013) is significantly better than Corridor 7, $t(5)=11.6$, $p=0.019$; Counter Strike 1.6, $t(5)=22.5$, $p=0.002$; and HROT, $t(5)=8.1$, $p=0.034$. In contrast to Wolfram, Doom, and Blitz Brigade—with an average $\Delta A$ of $19\%$, $18\%$, and $7\%$ respectively—the worst performing games are HROT, Corridor 7, and Counter Strike 1.6—with an average $\Delta A$ of $28\%$, $41\%$, and $50\%$, respectively across all LLMs in the $\Delta t=-2$ and $\theta=0.05$ setting.
Notably, in games which are easier to predict the proportion of the ground truth majority class is lower—Wolfram ($57\%$), Doom ($53\%$), and Blitz Brigade ($51\%$)—while games where the change in engagement is harder to predict have a higher naive baselineHROT ($61\%$), Corridor 7 ($71\%$), and Counter Strike 1.6 ($78\%$). This indicates that engagement prediction is easier in where the provided frames capture a more dynamic gameplay footage and have a more uniform distribution of increasing vs. decreasing engagement labels.
@@
Considering the overall performance of LLM engagement prediction across games, we fix our parameters for processing the ground truth at $\Delta t=-2s$ and $\theta=0.05$ for the remaining experiments presented in this paper.
V-C Engagement Prediction Based on Text Descriptions
Figure 6: Model performance on Text Descriptions across Basic and Advanced descriptions, and the best Multimodal One-Shot results identified in Section V-B. The table presents $\Delta A$ values (relative gain in accuracy). The last column show the average $\Delta A$ across all games. The last two rows show the MLP and CNN Baselines.
@@ V-C Engagement Prediction Based on Text Descriptions
In this section we examine the impact of text-based vs. multimodal prompting strategies on LLM performance. While in the former case we provide solely a text prompt to the model, in the latter case we feed both a text prompt and a corresponding image. Because the performance of LLMs can be affected even by small prompt variations [41], we experiment with both Basic and Advanced prompts. The prompting procedure for the text-based experiments are detailed in Section IV-D. Figure 5 shows the $\Delta A$ performance of Text Descriptions experiments compared to the best Multimodal One-Shot models discussed in the previous section.
@@
In this section our analysis focuses on the Text Descriptions compared to the Multimodal One-Shot results presented in the previous section. This focus on text allows us to compare the Text Descriptions method to a simple multimodal approach across different models. Our hypothesis is that the strategy of generating text-descriptions of frames first and then using these descriptions as part of the Text Descriptions will improve model performance, because it essentially encodes the images in terms of action and player involvement. We assume that using this type of Text Descriptions will present a better representation by discarding surface-level differences between frames and emphasising the structural differences.
We make four types of comparison: 1) we compare models within the Basic and Advanced Text Description tasks; 2) we compare between the downstream tasks of LLMs (Basic and Advanced Text Description, and Multimodal One-Shot) and the Baseline performance within each model; 3) we compare performance on individual games across all LLMs within the Basic and Advanced Text Description tasks; and finally 4) we compare performance on individual games between the the downstream tasks of LLMs (Basic and Advanced Text Description, and Multimodal One-Shot). In the first two cases, our degree of freedom is 19, while in the last two cases it is 5. We make 15, 17, 190, and 60 comparisons per test, respectively.
In general, the overall results of the experiments on Text Descriptions produce similar results to the Multimodal One-Shot setup. While we can observe that the overall performance is better on Advanced Text Descriptions than on Basic Text Descriptions; and that Multimodal One-Shot outperforms both—when we compare models both within tasks and between tasks, no significant results are observed. The same is true when comparing them to the MPL and CNN Baselines. However, several significant differences emerge when comparing individual game performance across all LLMs within the Basic and Advanced Text Description tasks. Most notably, Wolfram significantly outperforms PUBG, $t(5)=17.8$, $p=0.011$ and Counter Strike 1.6, $t(5)=14.6$, $p=0.015$; and Fortnite outperforms Counter Strike 1.6, $t(5)=11.5$, $p=0.032$ in the Advanced Text task. In the Basic Text task, Wolfram outperforms Medal of Honor, $t(5)=8.8$, $p=0.031$; Corridor 7, $t(5)=9.6$, $p=0.029$; Counter Strike 1.6, $t(5)=9.7$, $p=0.029$; and Borderlands 3, $t(5)=10.2$, $p=0.029$. Doom outperforms Heretic, $t(5)=11.1$, $p=0.028$; Counter Strike 1.6, $t(5)=11.2$, $p=0.028$; and Void Bastards, $t(5)=9.2$, $p=0.029$. Finally, both FarCry and Blitz Brigade outperforms Counter Strike 1.6, $t(5)=14.0$, $p=0.023$ and $t(5)=13.3$, $p=0.023$; and Corridor 7, $t(5)=9.1$, $p=0.029$ and $t(5)=9.5$, $p=0.029$ respectively. Finally, comparing task performance within individual games, we find that the Multimodal One-Shot task significantly outperforms Advanced Text Description on Counter Strike 1.6, $t(5)=5.0$, $p=0.023$; while both Basic and Advanced text tasks outperform Multimodal on Borderlands 3 (2019), $t(5)=3.5$, $p=0.047$ and $t(5)=4.0$, $p=0.047$ respectively.
Similarly to the Multimodal One-Shot experiments, Wolfram, Doom, and Blitz Brigade are the easiest games to predict; and Counter Strike 1.6, Heretic, and Corridor 7 are the hardest to predict. While there are some standout performances that we can observe on 5—for example Phi 3.5 is comparable to GPT-4o on Wolfram but performs poorly on Doom in the Advanced Text Description task compared to every other model—there is no apparent overarching pattern we can analyse. It also seems that any performance outliers can mostly be explained through the particularities of the data and the chosen algorithm.